OpenStudy (anonymous):

Write an equation of the line that passes through (8, 3) and (4, 3) A. y = 3 B. y = 4x + 3 C. y = 4 D. y = 3x

jimthompson5910 (jim_thompson5910):

The y coordinates are both points are 3, which means what?

OpenStudy (anonymous):

I think is A ;)

OpenStudy (campbell_st):

find the gradient of the line \[m = \frac{(3 -3)}{8-4)} = 0\] the gradient of the line is 0 so the line is horizontal and parallel to the x axis. using the point slope with either point y - 3 = 0(x - 4) y - 3 = 0 or y = 3